A sample of hydrazine sulphate (N2H6SO4) was dissolved in 100 mL of water 10 mL of this solution was reacted with excess of ferric chloride solution and warmed to complete the reaction. Ferrous ion formed was estimated and it, required 20 mL of M/50 potassium permanganate solution. Estimate the amount of hydrazine sulphate in one litre of the solution.
Reaction
4Fe3+ + N2H4 → N2 + 4Fe2+ + 4H+
〖MnO〗_4^- + 5Fe2+ + 8H+ → Mn2+ + 5Fe3+ + 4H2O

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol Meq of 〖MnO〗_4^- required = 20 × 1/50 × 5 = 2 ⇒ Meq of Fe2+ present in solution = 2 ⇒ millimol of Fe2+ present in solution = 2 (n-factor = 1) Also, ∵ 4 millimol of Fe2+ are formed from 1 millimol N2H4 ∴ 2 millimol Fe2+ from 1/4 ×2= 1/2 millimol N2H4 Therefore, molarity of hydrazine sulphate solution 1/2 × 1/10= 1/20 ⇒ In 1 L solution 1/20 mol N2H6SO4 is present. ⇒ Amount of N2H6SO4 = 1/20 × 130 = 6.5 gL-1